# zkSTAR: A zero knowledge system for time series attack detection enforcing regulatory compliance in critical infrastructure networks

### v1

@@ V-A Temporal Decomposition Scheme
Hypothesis Test Kernel: We conduct the $\chi^{2}$ hypothesis test using the SC-Kernel which considers the accumulated residual and covariance values at the end of every detection window. Additionally, we compute the SVD of the accumulated covariance matrix in an offline manner and validate its consistency using the SVD reconstruction kernel denoted by SVD-Kernel. More specifically, we represent the inputs to the SVD kernel in terms of $\mathcal{Z}^{SVD}_{W}$ as stated in Equation (19) which consists of the singular vectors and the square roots of the singular values.
@@
V-D zkSTAR Algorithmic Framework
We describe the overall zkSTAR algorithmic framework in Algorithm 6 that encapsulates the temporal and statistical consistency kernels. The framework primarily executes on a rolling horizon of detection windows characterized by $\mathcal{W}_{i}$ with each window consisting of $W$ TC intervals of $D$ timestamps each.
@@
Algorithm 6 zkSTAR Algorithmic Framework
1:#Set visibility for TC circuit
[…]
3:
4:#Setup proving circuit for TC
[…]
6:
7:#Set visibility for SC circuit
[…]
9:
10:#Setup proving circuit for SC
[…]
12:
13:distribute $vk^{TC},vk^{SC}$ publicly
14:
15:for $i=1,2\ldots$ do $\triangleright$ detection window $\mathcal{W}_{i}$
16:for $j=1,2,\ldots W$ do $\triangleright$ TC interval $w_{ij}$
17: generate nonce values $\mathcal{N}^{TC}_{j}$
18: compute $\mathcal{Z}^{TC}_{j}$ using Equation (17)
19: compute $\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1}$ using Equations (13)-(15).
[…]
22:
23:#Generate TC proof for interval $j$
[…]
25:
26: extract $\mathcal{X},\mathcal{R},\mathcal{K}]_{j}$ from $\mathcal{O}^{TC}_{j}$
27:end for
28: extract $S_{W}$ from $\mathcal{R}_{W}$
[…]
30: generate nonce values $\mathcal{N}^{SVD}$
31: compute $\mathcal{Z}^{SVD}_{W}$ using SVD values $U_{W},\Sigma_{W}^{-\frac{1}{2}},\mathcal{N}^{SVD}$
[…]
33:
34:#Generate SC proof for window $\mathcal{W}_{i}$
[…]
36:
37:return $\mathcal{P_{W_{i}}}=\Big[\Big(\Pi^{TC}_{w_{i1}}\ldots\Pi^{TC}_{w_{iW}}\Big),\Pi^{SC}_{\mathcal{W_{i}}}\Big]$
38:end for
For each detection window $\mathcal{W}i$, the algorithm iterates through the TC intervals wijw{ij} of length $W$. Within each TC interval, a fresh set of nonce values $\mathcal{N}^{TC}{j}$ is generated to enable efficient hiding for hash outputs. Using these nonces, the algorithm constructs the nonce-appended input set $\mathcal{Z}^{TC}{j}$ according to (17), which encapsulates sensor data, model Jacobians, and Kalman gain parameters. The posterior state, residual, and kernel reconstruction variables from the previous interval, $\mathcal{X},\mathcal{R},\mathcal{K}]{j-1}$, are then obtained from (13)–(15). Together, these form the TC-kernel input $\mathcal{I}^{TC}{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]{j-1},[\mathcal{Z}^{TC},\mathcal{N}^{TC}]{j}}$, which is passed to the TC-Kernel for zkSNARK-based verification of temporal consistency. The kernel outputs the updated tuple $\mathcal{O}^{TC}{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]{j}}$, whose elements are extracted and propagated to the next interval.
@@
VI Experimental Results
@@
Figure 5: HAI Dataset: Detection Quality with varying PSF and fixed D
For experimental studies, we utilize the HAI [42] and ORNL-PS [20, 43] dataset. To build a asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section III. We encode the temporal and statistical components of these models in corresponding zk-SNARKs as described in previous sections. As a result, the zkSTAR framework yields that encode temporal and statistical consistency with respect to detection outcomes.
We carry out detection based on a pre-determined value of upper control limit (UCL) that is representative of the levels of significance used for the $\chi 2$ hypothesis test. These UCL values are generally dependent on the subsystem data and might also rely on engineering domain expertise as well. We use a value of 5e35e-3 and 2e22e-2 for HAI and ORNL-PS datasets that were obtained after rigorous empirical studies.
System Implementation Details: All experiments were carried out on a virtual machine (VM) running Ubuntu 24.04 with 100GB of RAM and 16 vCPUs using Python 3.11 with the detection model inference using PyTorch 2.7.1. We utilize the ezkl library for generating all the zkSNARKs associated with the zkSTAR frameworkWe provide a native as well as a container based mechanism for executing the zkSTAR framework. The native method is primarily meant for comprehensive analysis on a high-performance computing (HPC) clusters to evaluate detection accuracy and computational correctness with varying value of residual aggregation window (referred to as the D value) and precision scale factor (PSF value). The container based method is primarily meant to evaluate real-world system performance in terms of CPU utilization and memory consumption for the distinct processes of witness generation as well as temporal consistency and statistical consistency proofs. We utilize separate containers to represent utility as the prover and regulator as the verifier with an interactive REST API driven interface for evaluation and deployment.
VI-A Detection Quality of zkSNARK outcomes
For assessing the detection quality, we carry out two sets of experiments that depict the effects of varying the PSF with a fixed D and vice versa as well. Specifically, these experiments pertain to a total detection run-time window of 10 minutes in which the attack starts roughly in the middle. Additionally, the attack duration window includes the actual attack occurrence as well as a buffer time of 3 minutes for the HAI dataset. For the ORNL-PS dataset, the attack duration window lasts until the end of the detection run-time window of 10 minutes. Even though the HAI attack scenario lasts 164 secs, we include a buffer time for HAI due to the highly non-linear nature of the detection model that can induce slight delays before anomalies can acquire statistical significance. For these experiments, we consider PSF values of 8,10 and 12, while using 4, 8, 16 and 32 as the D values.
@@ VI-A Detection Quality of zkSNARK outcomes
In Figures 5 and 6, we depict the detection quality with respect to varying PSF values for distinct values of D. In this case we can observe that the lower PSF values usually exhibit greater volatility in detection, while higher PSF values correspond to a less volatile nature of the test statistic. Additionally, we also observe a slight increase in volatility with increasing D which is expected since the errors in residual computation compound for lower PSF values with higher D values.
From these experiments, we can conclude that a higher PSF value might be more desirable for a lower D. However, a mid-range PSF value might also suffice for a higher D value. Regardless, the experiments establish the robustness in detection quality of the zkSTAR framework. The setup times for HAI and ORNL-PS dataset has been provided in Tables III and IV respectively.
@@
Table I: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for Temporal Consistency (mean(std) in MB)
@@
Table II: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for Statistical consistency (mean(std) in MB)
Setup Setup Time (s)
Phase D=1 D=4 D=8 D=16
Model Export 24.77 3.25 7.96 24.21
Settings Generation 30.02 3.53 9.55 28.99
Circuit Compilation 24.46 2.24 7.06 24.90
Circuit Setup 171.63 146.84 162.04 167.55
Table III: HAI Setup Time Breakdown (in seconds)
Setup Setup Time (s)
Phase D=1 D=4 D=8 D=16
Model Export 24.77 3.25 7.96 24.21
Settings Generation 30.02 3.53 9.55 28.99
Circuit Compilation 24.46 2.24 7.06 24.90
Circuit Setup 171.63 146.84 162.04 167.55
Table IV: ORNL-PS Setup Time Breakdown (in seconds)
@@ D HAI ORNL-PS
Table V: Temporal Consistency Witness Size with Varying D (mean(std) in MB)
@@
8 56.06(22.60) 90.83(5.41) 1.00 80.81(4.19) 4.50 70.61(2.87) 10.29 78.25(3.75 18.57
10 54.26(19.17) 76.16(1.55) 1.00 71.32(3.22) 4.27 75.97(8.88) 8.02 77.11(5.01) 15.80
12 53.49(6.60) 76.74(3.05) 1.00 66.50(2.34) 4.62 74.59(3.19) 8.23 77.79(3.90) 15.78
Table VI: HAI Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
@@
Table VII: ORNL-PS Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
VI-B Computational Correctness of zkSNARK outputs
Figure 7 denotes box plot series of relative error for varying D values for HAI and ORNL-PS datasets with PSF=12 pertaining to the SVD of the predicted covariance matrices. We note that in case of both datasets, the relative error remains well below 1 for a majority of the cases. This indicates that the reconstruction error remains less than 0.1 in most cases. We observe higher incidence of relative error values going above 1 during the attack window which can be due to the greater volatility in residual values computed during an attack. However, even in the worst case, the relative error with respect to matrix elements does not exceed 1.4 and 1.2 respectively for HAI and ORNL datasets.
Figure 8 denotes scatter plot series of relative errors for varying D values. We note that unlike the SVD check, the Kalman gain check is carried out after each D time steps, resulting in a staggered trend for the first few iterations. The Kalman gain check for both datasets conclusively demonstrates a consistently stable and low error value in the neighborhood of 1e51e-5 for HAI dataset. However, for the ORNL-PS dataset, we see a higher but highly stable relative error values aroun 1e21e-2.
VI-C System Performance of Utility Prover
Analysis of Proof and Key Sizes: Tables I and II presents the sizes of the proof, proving and verification keys for the temporal consistency and statistical consistency proofs. Table II also presents the witness size pertaining to the statistical consistency proofs. Additionally, we present the witness and the compiled circuit size for temporal consistency proofs separately in Table V.
In Table I, we can observe that for the case of temporal consistency, the proof size as well as proving and verification key sizes are similar for both datasets. On the other hand, in Table II, the circuit, proving and verification key size for the statistical consistency proofs were constant as expected and minor deviations are observed for the proof and witness sizes across both datasets.
However, we observe an interesting trend with respect to witness and circuit sizes as represented in Table V. It can be seen that the circuit and witness sizes for both datasets exhibit a drop from D values of 1 to 4 before gradually increasing for more extreme D values such as 8, 16 and 32. This behavior can be potentially explained on the basis of row domain padding used by Halo2 proving systems [44] (such as ezkl) which pads rows to the nearest value of $2^{k}$. This behavior is also evidenced by the fact that the witness and circuit sizes almost double from 4 to 12. As a consequence, we can conclude that D=4 might correspond to a sweet spot for the non linear Kalman Filter LSTM based circuit leading to a lower circuit and witness size.
Analyses of Circuit Setup Costs: In Tables III and IV, we present the time incurred during the setup phase of zkSTAR pertaining to model export, settings generation, circuit compilation and circuit setup for HAI and ORNL-PS datasets. In each of these tables we see an almost identical trend as those seen in Table V for each category of setup steps. These trends can be explained on the basis of the fact that the setup of zk-SNARKs relies significantly on the size of the circuits themselves. Therefore, the trends in Tables III and IV can be a natural consequence of the row domain padding adopted by Halo 2 proving backends. This ultimately results in trends wherein a decrease in time consumption is observed from D value of 1 to 4 followed by an increase for more extreme values like 8 and 16.
@@ VI-C System Performance of Utility Prover
Computational Analysis of Proving System: We analyze the computational performance for temporal and statistical consistency proof generation for both datasets for varying D values. Our results are presented in Tables VI and VII for HAI and ORNL-PS datasets respectively. The data for these experiments are obtained from multiple runs of varying PSF values and represent the mean and standard deviation for one proof invocation of the circuit. We also provide a speed up value that represents the average case speedup that can be expected when sequentially generating proofs for the entire horizon. We provide the speedup formula and its associated explanation in Equation (58) and Appendix A-A respectively.
@@
The temporal consistency proof generation times are represented by the mean and standard deviation values for D values of 1,4,8 and 16. The proving times are in general stable with minor variations observed with varying D values. We observe that the mean proof generation time decreases slightly from D values of 1 to 8, before increasing slightly at 16. This trend is consistently observed across both datasets for all PSF values. The non-monotonous trends depicted in Tables VI and VII can be due to the non-linear proving costs afforded by ezkl backend. Fundamentally, this trend can be explained on the basis of Table V which highlights the non monotonous behavior of witness and circuit sizes resulting from row domain padding used by Halo2 proving systems. Since the proving times are directly proportional to the circuit size, the trends in Tables VI and VII are exactly aligned with those depicted in Table V.
@@
Lastly, the hypothesis test columns in Tables VI and VII reflect a relatively stable values of average proving time corresponding to statistical consistency proofs. These values are slightly higher for HAI compared to ORNL-PS which might be explained on the basis of the difference in dimensionality between both problems.
@@
System Performance of Witness and Proof Generation: Figures 9 and 10 present trends pertaining to the CPU utilization and memory consumption for only the HAI dataset. we provide the corresponding trends for ORNL-PS in Appendix A-B due to space constraints as well as similar nature of both curves. From Figure 9, we can see greater volatility in CPU utilization and memory consumption for D value of 1. We notice that this trend gradually stabilizes for D values of 8 and 16. The underlying reason for this trend is that proof generation is an inherently lower values of D demand more frequent proving steps that results in higher volatility in CPU utilization and memory consumption. As D grows to 8 and 16, the proving time stabilizes on account of less frequent proving steps encountered during the simulation horizon.
Additionally, we present the proving time system performance using CPU utilization and memory consumption for a random proving window for varying D values. We observe a steady uptick in CPU utilization and memory consumption reaching a value of 1600% and 55GB respectively. The CPU utilization trends reflect the multi-processing capabilities on the hardware that was used which comprised of 16 vCPUs. On the other hand, the memory consumption of 55GB indicates that the proving process consumed more than half the total memory available on the device.
VII Conclusion

### v2

@@ 5.1 Temporal Decomposition Scheme
Hypothesis Test Kernel: We conduct the $\chi^{2}$ hypothesis test using the SC-Kernel which considers the accumulated residual and covariance values at the end of every detection window. Additionally, we compute the SVD of the accumulated covariance matrix in an offline manner and validate its consistency using the SVD reconstruction kernel denoted by SVD-Kernel. We represent the inputs to the SVD kernel in terms of $\mathcal{Z}^{SVD}_{W}$ as stated in Equation (20) which consists of the singular vectors and the square roots of the singular values.
@@
Algorithm 6 zkSTAR Algorithmic Framework
1:#Set visibility for TC circuit
2:#input;output;param visibility
[…]
4:
5:#Setup proving circuit for TC
[…]
7:
8:#Set visibility for SC circuit
9:#input;output;param visibility
[…]
11:
12:#Setup proving circuit for SC
[…]
14:
15:distribute $vk^{TC},vk^{SC}$ publicly
16:
17:for $i=1,2\ldots$ do $\triangleright$ rolling detection windows $\mathcal{W}_{i}$
18:for $j=1,2,\ldots W$ do $\triangleright$ TC interval $w_{ij}$
19: generate nonce values $\mathcal{N}^{TC}_{j}$
20: compute $\mathcal{Z}^{TC}_{j}$ using Equation (18)
21: compute $\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1}$ using Equations (14)-(16).
[…]
24:
25:#Generate TC proof for interval $j$
[…]
27:
28: extract $\mathcal{X},\mathcal{R},\mathcal{K}]_{j}$ from $\mathcal{O}^{TC}_{j}$
29:end for
30: extract $S_{W}$ from $\mathcal{R}_{W}$
[…]
32: generate nonce values $\mathcal{N}^{SVD}$
33: compute $\mathcal{Z}^{SVD}_{W}$ using SVD values $U_{W},\Sigma_{W}^{-\frac{1}{2}},\mathcal{N}^{SVD}$
[…]
35:
36:#Generate SC proof for window $\mathcal{W}_{i}$
[…]
38:
39:return $\mathcal{P_{W_{i}}}=\Big[\Big(\Pi^{TC}_{w_{i1}}\ldots\Pi^{TC}_{w_{iW}}\Big),\Pi^{SC}_{\mathcal{W_{i}}}\Big]$
40:end for
5.4 zkSTAR Algorithmic Framework
We describe the overall zkSTAR algorithmic framework in Algorithm 6 that encapsulates the TC and SC kernels. The framework primarily executes on a rolling horizon of detection windows characterized by $\mathcal{W}_{i}$ with each window consisting of $W$ TC intervals of $D$ timestamps each.
@@
For each detection window $\mathcal{W}i$, the algorithm iterates through the TC intervals $w_{ij}$ of length $W$. Within each TC interval, a fresh set of nonce values $\mathcal{N}^{TC}_{j}$ is generated to enable efficient hiding for hash outputs. Using these nonces, the algorithm constructs the nonce-appended input set $\mathcal{Z}^{TC}_{j}$ according to Equation (18), which encapsulates sensor data, model Jacobians, and Kalman gain parameters. The posterior state, residual, and kernel reconstruction variables from the previous interval, $\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1}$, are then obtained from (14)–(16). Together, these form the TC-kernel input $\mathcal{I}^{TC}_{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1},[\mathcal{Z}^{TC},\mathcal{N}^{TC}]_{j}}$, which is passed to the TC-Kernel for zkSNARK-based verification of temporal consistency. The kernel outputs the updated tuple $\mathcal{O}^{TC}{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]{j}}$, whose elements are extracted and propagated to the next interval.
@@
5.5 Assumptions and Limitations
The zkSTAR framework relies on several assumptions pertaining to the properties of the zero-knowledge proving system. First, since zkSTAR uses the Halo2 proof generation paradigm as a consequence of using the ezkl framework [ezkl2024], it implicitly depends on elliptic-curve cryptographic primitives and on the integrity of the Common Reference String (CRS). As a result, we assume that the CRS is generated through a publicly verifiable ceremony per-utility. Second, we assume that the EKF model is well-formed and is strongly sensitive to the Jacobian matrices [huang2008analysis] implying that slight deviations in SVD or Kalman gain computations would lead to inconsistent state and residual estimates violating either TC or SC conditions. Third, we assume that verifier is honest-but-curious and that public values of the proof artifacts do not violate privacy of utility state space which is a common requirement of existing regulators [doe]. Finally, we assume that communication channels are assumed to preserve integrity but not confidentiality which can be alleviated by using transport-level encryption like TLS for privacy during proof transmission. Under these assumptions, zkSTAR achieves end-to-end regulatory verification with cryptographic soundness and provable privacy, while maintaining computational feasibility for real-world industrial deployments.
6 Experimental Results
@@
For experimental studies, we utilize the HAI [hai] and ORNL-PS [pan2015classification, pan2015developing] datasets. To build an asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section 3. The training methodology for EKF models is briefly discussed in Appendix A.2. We encode the temporal and statistical components of these models in corresponding zk-SNARKs as described in previous sections. As a result, the zkSTAR framework encodes temporal and statistical consistency with respect to detection outcomes.
We carry out detection based on a pre-determined value of upper control limit (UCL) that is representative of the levels of significance used for the $\chi 2$ hypothesis test. These UCL values are generally dependent on the subsystem data and might also rely on engineering domain expertise as well. We use a value of 5e35e-3 and 2e22e-2 for HAI and ORNL-PS datasets that were obtained after rigorous empirical studies. All of our results include comparison with the non-ZKP version as our benchmark to judge the detection quality.
System Implementation Details: All experiments were carried out on a virtual machine (VM) running Ubuntu 24.04 with 100GB of RAM and 16 vCPUs using Python 3.11 with the detection model inference using PyTorch 2.7.1. We utilize the ezkl library [ezkl2024] for generating all the zkSNARKs associated with the zkSTAR framework. We implement zkSTAR using both native and containerized execution modes. The native mode targets high-performance computing (HPC) clusters for evaluating detection accuracy and computational correctness across varying residual aggregation windows ($D$) and precision scale factors (PSF). The containerized mode assesses real-world performance, including CPU and memory usage during witness generation and TC and SC proof generation and verification. Separate containers emulate the utility (prover) and regulator (verifier), communicating via REST APIs and managed through the zkstarctl command-line interface.
6.1 Detection Quality of zkSNARK outcomes
To evaluate detection quality, we conduct two experiments examining the effects of varying the PSF with fixed $D$ and vice versa. For brevity, PSF variation results are shown in Appendix C.2. Each experiment spans a 10-minute detection window, with attacks initiated mid-run. For HAI, the attack window includes the actual 164s event and a 3-minute buffer to account for the non-linear model’s delayed anomaly convergence, whereas in ORNL-PS, the attack persists until the end of the run. We test PSF values of 8, 10, and 12, and $D$ values of 4, 8, 16, and 32.
@@ 6.1 Detection Quality of zkSNARK outcomes
6.2 Computational Correctness of zkSTAR output
Figure 5 shows box plots of relative error for varying D values on the HAI and ORNL-PS datasets (PSF = 12) based on the SVD of predicted covariance matrices. For both datasets, relative errors remain below 1 in most cases, indicating reconstruction errors under 0.1. Slight error spikes above 1 occur during attack windows due to increased residual volatility, yet even in the worst case, relative errors stay below 1.4 (HAI) and 1.2 (ORNL-PS). Figure 6 shows scatter plots of relative errors across varying $D$ values. Unlike the SVD check, the Kalman gain check, which is performed every $D$ steps, exhibits an initial staggered trend. For the HAI dataset, errors remain consistently low around 1e51e-5, while for the ORNL-PS dataset they are slightly higher yet stable, around 1e21e-2.
6.3 System Performance of Utility Prover
@@
Table I: TC Circuit: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for TC (mean(std) in MB)
@@
Table II: SC Circuit Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for SC in MB)
@@ D HAI ORNL-PS
Table III: Temporal Consistency Witness Size with Varying D (mean(std) in MB)
Analysis of Proof and Key Sizes: Tables I and II presents the sizes of the proof, proving and verification keys for the temporal consistency and statistical consistency proofs. Table II also presents the witness size pertaining to the statistical consistency proofs. Additionally, we present the witness and the compiled circuit size for temporal consistency proofs for PSF value 8 separately in Table III.
In Table I, we can observe that for the case of temporal consistency, the proof size as well as proving and verification key sizes are similar for both datasets. On the other hand, in Table II, the circuit, proving and verification key size for the statistical consistency proofs were constant as expected and minor deviations are observed for the proof and witness sizes across both datasets. However, we observe an interesting trend with respect to witness and circuit sizes as represented in Table III. Mean witness sizes increase gradually from around 43kB to 360kB for HAI and 36kB to 325kB for ORNL-PS. On the other hand, the circuit sizes increase from a smaller value of 0.25MB at D value of 1 to 5.8MB for HAI while exhibiting a near identical trend for ORNL-PS dataset as well. This trend is natural and expected since increase in D value corresponds to larger TC interval sizes making the TC circuit grow exponentially.
@@
8 56.06(22.60) 70.05(3.03) 1.00 80.81(4.19) 3.5 70.61(2.87) 7.93 78.25(3.75) 14.32
10 54.26(19.17) 69.26(3.22) 1.00 71.32(3.22) 3.89 75.97(8.88) 7.3 77.11(5.01) 14.37
12 53.49(6.60) 69.89(3.77) 1.00 66.50(2.34) 4.20 74.59(3.19) 7.5 77.79(3.90) 14.37
Table IV: HAI Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
@@
Table V: ORNL-PS Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
@@
Computational Analysis of Proving System: We analyze the computational performance for temporal and statistical consistency proof generation for both datasets for varying D values. Circuit generation and setup costs are presented separately in Appendix C.4. Our results are presented in Tables IV and V for HAI and ORNL-PS datasets respectively. The data for these experiments are obtained from multiple runs of varying PSF values and represent the mean and standard deviation for one proof invocation of the circuit. We also provide a speed up value that represents the average case speedup that can be expected when sequentially generating proofs for the entire horizon. We provide the speedup formula and its associated explanation using Equation (48) in Appendix C.1 respectively.
@@
The temporal consistency proof generation times are represented by the mean and standard deviation values for D values of 1,4,8 and 16. The proving times are in general stable with minor variations observed with varying D values. We observe that in case of HAI, the mean proof generation time increases slightly going from D values of 1 to 4 before decreasing slightly from 4 to 8 and finally increasing from 8 to 16. On the other hand the trend is inverted for ORNL-PS, decreasing slightly from 1 to 4, increasing from 4 to 8 and then decreasing again from 8 to 16. Interestingly, there is more volatility for PSF value of 8 in case of either dataset as compared to 10 and 12. The non-monotonous trends depicted in Tables IV and V can be due to the non-linear proving costs afforded by ezkl backend which could be explained on the basis of row domain padding used by Halo2 proving systems. Further, analyzing speedups, we note a healthy increase between each set of D values regardless of the PSF value or the dataset used. Since the circuit size is known to influence the proving times, the trends in Tables IV and V are exactly aligned with those depicted in Table III. The speedup value insights point to the fact that having a higher TC interval size (or D value) is almost always more computationally efficient than invoking proof generation on a finer granularity corresponding to a lower TC interval size.
@@
Lastly, the hypothesis test columns in Tables IV and V reflect a relatively stable values of average proving time corresponding to statistical consistency proofs. These values are slightly higher for HAI compared to ORNL-PS which might be explained on the basis of the difference in dimensionality between both problems.
@@
System Performance of Witness and Proof Generation: Figures 7 and 8 present trends pertaining to the CPU utilization and memory consumption for only the HAI dataset. We provide the corresponding trends for ORNL-PS in Appendix C.3 due to space constraints. The results show a steady and predictable scaling trend, with CPU utilization peaking around 1600%, demonstrating effective parallelization, and memory usage reaching approximately 55 GB, well within available system capacity. From Figure 7, we can see greater volatility in CPU utilization and memory consumption for D value of 1 eventually stabilizing for D values of 8 and 16. The underlying reason for this trend is that lower values of D demand more frequent proving steps that results in higher volatility in CPU utilization and memory consumption. As D grows to 8 and 16, the proving time stabilizes on account of less frequent proving steps encountered during the simulation horizon. These results confirm that zkSTAR efficiently leverages multi-core resources with stability in memory consumption under increasing computational load.
Verification Latency: To gauge the regulatory overhead in terms of verification, we measure the verification times incurred by the regulator for verifying TC and SC proof artifacts. In almost all cases, zkSTAR SC and TC proofs are successfully verified in under 1.5s for both datasets with varying TC interval sizes. Detailed results pertaining to verification times for each case has been presented in Appendix C.5. These results assert the practicality of zkSTAR for near real-time regulatory audits with minimal overhead.
7 Conclusion